Supplemental Operations Report – October 15th

Date of supplemental Operations Report: 10/15/2023
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Bread Maker – inside bucket is leaking and needs to be replaced, replacement is already in order. New toilet tank lever sensor shows symptoms of malfunctioning or being covered with toilet paper, further diagnostic required.

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 48 %; Avg 69.3%.
VDC Last 24 hours: Max 58.44V; Min 43.81V; Avg 49.12V.
Generator run time: Not used yet, serviced during work party. Hours – 3827.8.

Propane Readings:
Station Tank: 72%
Director Tank: 50%
Intern Tank: 66%
Generator Tank: 76%

Hab Static Tank – 500 gallons
GreenHab – 200 gallons
Outpost tank – 250 gallons
Science Dome – 0 gallons
Hab Toilet Tank emptied: Yes

Sojourner rover used: Yes. From Hanksville to the station and around MDRS
Hours: 195.2
Beginning Charge: 100 %
Ending Charge: 37 %
Currently Charging: Yes
Notes on Rovers: All rovers refilled. Perseverance – got seats replaced, long battery wire needs to be replaced some day in the future. Curiosity – two short battery wires replaced. Spirit – one short battery wire replaced. Opportunity – two short battery wires need to be replaced.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: To Hanksville for supplies. To Grand Junction for supplies.
General notes and comments: Hab Car – need to be serviced. Crew Car – Serviced. Engine oil – nominal, engine coolant – refilled, tires – rotated.

Summary of Internet: Nominal.

EVA suits and radios: Suits: All suits are serviced and ready for the season.
Comms: Radio repeater setted up on the North Ridge. All EVA radios are ready for the season.

Campus wide inspection, if action taken, what and why: Campus cleaned and ready for the season. Septic tank emptied.
Summary of Hab Operations: All systems are ready for the season. Toilet tank has been resealed and sensors have been added to monitor its level.
Summary of GreenHab Operations: Cleaned and ready to be planted. All operations scheduled to be done during crew 283 rotation. Old and damaged watering cans replaced.
Summary of SciDome Operations: Cleaned and organized for the season. iMac was setted up for the EVA Link project. Smart Home server plugged in.
Summary of Observatories Operations:

Cleaned and serviced by Astronomy crew.
Summary of RAM Operations: Cleaned and restocked for the season.
Summary of Outpost Operations: Cleaned and ready for the season. Septic tank emptied. The Director’s trailer heater went offline due to a mouse going inside. Inspection revealed there could be a problem with the control board. Replacement of the board is in order.
Summary of Health and Safety Issues: Nothing to report. Standart trauma kit ordered for EVA.

Copyright © The Mars Society. All rights reserved. | Main Site